A customer relations manager at observes that a teller can serve, on average, 15 clients in an hour. Calculate the probability that the teller serves: 6 clients in a 30 minute period. Interpret your answer. More than 14 but less than 17 clients in a particular hour. Interpret your answer.


1
Expert's answer
2021-09-27T23:29:00-0400

Let "X=" the number of clients served by a teller: "X\sim Po(\lambda t)."

a) "\lambda=15, t=0.5"


"\lambda t=15(0.5)=7.5"

"P(X=6)=\dfrac{e^{-\lambda t}(\lambda t)^6}{6!}=\dfrac{e^{-7.5}(7.5)^6}{6!}"

"\approx0.136718"

The probability that the teller serves 6 clients in a 30 minute period is "0.136718."


b) "\lambda=15, t=1"

"\lambda t=15(1)=15"

"P(14<X<17)=P(X=15)+P(X=16)"

"=\dfrac{e^{-15}(15)^{15}}{15!}+\dfrac{e^{-15}(15)^{16}}{16!}"

"\approx0.1024359+0.0960336\approx0.198470"


The probability that the teller serves more than 14 but less than 17 clients in a particular hour is "0.198470."